fing the missing term in the arithemetic sequence
55,_,_,_,155

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

find the missing term in the arithemetic sequence
55,_,_,_,155
-----
a(5) = a(1) + 4d
======
155 = 55 + 4d
4d = 100
d = 25
------
Sequence::
55, 80, 105, 130, 155
=========================
Cheers,
Stan H.
=======================

Answer by richwmiller(17219)

a=55
a+4(d)=155
4(d)=100
d=25
55, 80, 105, 130, 155